espress the area for a rectangle with a perimeter of 100ft as a function of its length
Answer by stanbon(75887)

espress the area for a rectangle with a perimeter of 100ft as a function of its length
-------------------
P = 2(L + W)
---
100 = 2(L+W)
L + W = 50
W = 50-L
----------------
Area = (length*width)
---
Area = L(50-L)
Area = 50L - L^2
====================
Cheers,
Stan H.
================